OpenStudy (anonymous):

Solve the equation: t/3 - 1/2 = 1

OpenStudy (anonymous):

add one half, multiply by three

OpenStudy (anonymous):

so 1/2 + 1/2 and add the three where?

OpenStudy (anonymous):

t/3=1

OpenStudy (anonymous):

t= 9/2

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

\[1+\frac{1}{2}=\frac{3}{2}\] \[3\times \frac{3}{2}=\frac{9}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is it as a decimal

OpenStudy (anonymous):

is it one of theses 4.1 B. 5.0 C. 4.2 D. 4.5

OpenStudy (anonymous):

Hello?????

OpenStudy (anonymous):

\[\frac{ t }{ 3 }-\frac{ 1 }{ 2 }=1\] Next, you reverse addition with subtraction; \[\frac{ t }{ 3 }-\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }=1+\frac{ 1 }{ 2 }\] \[\frac{ 1 }{ 2 }=0.5\] \[\frac{ t }{ 3 }=1.5\] Next, you reverse division with multiplication; \[\frac{ t }{ 3 }\times3=1.5\times3\] \[1.5\times3=4.5\] \[t=4.5\] So your answer is D. 4.5

OpenStudy (anonymous):

Meant to say reverse subtraction with addition, my bad. :p

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

No problem. (: